OpenStudy (anonymous):

Help ASAP!! Will FAN & MEDAL :) Lance has 38 eggs that he needs to put into egg crates. The egg crates hold 12 eggs each. To calculate how many egg crates he needs to fit all his eggs, he divides 38 by 12 to get 3 egg crates. Did Lance do the problem correctly? No, because 38 divided by 12 equals 3 with remainder 2; therefore, 2 eggs will be left out No, because 41 divided by 12 equals 3 with remainder 5; therefore, 5 eggs will be left out Yes, because 38 divided by 12 equals 3 No, because 38 divided by 12 equals 4; therefore, he will need 4 crates

OpenStudy (anonymous):

I'm going with the 3rd one

OpenStudy (anonymous):

wait no i it's a

OpenStudy (triciaal):

this seem to be a trick question. 38/12 = 3 crates with 2 eggs needing 1 more crate so he will need 4 crates find multiples of 12 that is more than 38 so will be 4

OpenStudy (anonymous):

Well if we know we have a total of 32 eggs and each egg crate holds 12 eggs/crate then if we do what Lance did our units should be in "crate" Let us see if this is true \[\frac{ 38 eggs}{ \frac{ 12eggs }{ 1 crate }} = 38eggs \times \frac{ 1 crate }{ 12 eggs} = 3 crates\] Since the our value is 3 and our units are also correct, you are correct in that it would be the 3rd choice

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

My pleasure